A solar collector is placed in direct sunlight where it absorbs energy at the rate of 892 J/s for each square meter of its surface. The emissivity of the solar collector is

e = 0.77. What equilibrium temperature does the collector reach? Assume that the only energy loss is due to the emission of radiation.
Answer #1

According to Stefan-Boltzmann Law, The energy per unit time per unit area \small j is;

\small j=e\sigma T^{4}

Here \small j= 892 J/s for each square meter, e = 0.77 emissivity, \small \sigma = 5.67 \small \times 10-8 Wm-2K-4 is Stefan-Boltzmann constant and T is temperature. Thus;

  \small \Rightarrow T=\sqrt[4]{\frac{j}{e\sigma }}

  \small \Rightarrow T=\sqrt[4]{\frac{892}{0.77\times 5.67\times 10^{-8}}} =378\: K